# Approximate Quadratization of High-Order Hamiltonians for Combinatorial Quantum Optimization

### v1

@@ Abstract
Combinatorial optimization problems have wide-ranging applications in industry and academia. Quantum computers may help solve them by sampling from carefully prepared Ansatz quantum circuits. However, current quantum computers are limited by their qubit count, connectivity, and noise. This is particularly restrictive when considering optimization problems beyond the quadratic order. Here, we introduce Ansatze based on an approximate quadratization of high-order Hamiltonians which do not incur a qubit overhead. The price paid is a loss in the quality of the noiseless solution. Crucially, this approximation yields shallower Ansatze which are more robust to noise than the standard QAOA one. We show this through simulations of systems of 8 to 16 qubits with variable noise strengths. Furthermore, we also propose a noise-aware Ansatz design method for quadratic optimization problems. This method implements only part of the corresponding Hamiltonian by limiting the number of layers of SWAP gates in the Ansatz. We find that for both problem types, under noise, our approximate implementation of the full problem structure can significantly enhance the solution quality. Our work opens a path to enhance the solution quality that approximate quantum optimization achieves on noisy hardware.
@@ I Introduction
Quantum computing offers a new information-processing framework based on the laws of quantum physics. For certain tasks, quantum computing achieves better performance than the best known classical methods. Famously, Shor’s algorithm delivers an exponential speed-up in factoring over the best known classical algorithms [1]. However, this remains a theoretical feat since Shor’s algorithm requires deep circuits which cannot be faithfully implemented at practically relevant scales on current hardware due to noise. This has motivated the development of algorithms requiring shallow circuits such as Variational Quantum Algorithms (VQAs) [2]. VQAs are heuristic in nature and are often designed to find the ground state of a Hamiltonian [3, 4, 5]. VQAs optimize a parametrized quantum circuit called the Ansatz using a classical optimization procedure to update the parameters.
@@
The Ansatz in a VQA can be either problem-agnostic [6] or retain the structure of the problem to solve [7]. Problem-agnostic Ansatze are typically designed to minimize the effect of noise by accounting for hardware constraints. However, they come at the cost of limited expressivity compared to Ansatze that fully encode the problem structure. Designing expressive and shallow Ansatze remains an open question [8].
@@
Here, we focus on VQAs to solve combinatorial optimization problems, which are prevalent across industry applications such as communication [9], finance [10, 11], and vehicle routing [12, 13, 14]. Many relevant combinatorial optimization problems are hard to solve exactly [15]. In addition, some of them also do not have polynomial-time approximate algorithms that reach an arbitrarily accurate approximation of the optimal solution [16]. Therefore, many classical algorithms designed to solve such problems are heuristic in nature. Crucially, they work well in practice, thus motivating a similar heuristic approach for quantum computing [17].
@@
A foundational VQA for combinatorial optimization problems is the Quantum Approximate Optimization Algorithm (QAOA) [18, 19]. Its trotterized Ansatz structure alternates $p$ time evolutions of the target problem and of a problem-independent Hamiltonian. The structure is inspired from the adiabatic theorem in physics, and thus inherits its convergence guarantee in the limit $p\to\infty$ [20, 18, 21].
@@
Finite-depth QAOA is a heuristic algorithm, which implies that its performance must be thoroughly investigated and benchmarked on a case-by-case basis [17, 22]. In particular, research often investigates QAOA for quadratic unconstrained binary optimization problems (QUBOs), such as Max-Cut, to benchmark hardware performance [23, 24, 25, 26] and validate new algorithms [27, 28]. However, Max-Cut can often easily be exactly solved classically at low graph densities and up to 10 000 nodes [29]. Solving it using QAOA would only yield approximate solutions even in an ideal case [30, 31, 32, 33], thus making it a poor candidate to deliver a quantum advantage. Hence, recent studies investigated high-order unconstrained binary optimization problems (HUBOs) with cubic and quartic terms. Certain HUBOs can be very hard to solve, even at a low number of decision variables [34]. For instance, the largest exactly solved instance of the Low Autocorrelation Binary Sequence (LABS) problem has only $66$ decision variables [35, 36]. Importantly, Ref. [37] suggests a scaling advantage for QAOA over Gurobi [38] or CPLEX [39] in solving LABS.
@@
In this work, we first show in Section II how the quantum resources requirements of QAOA scale unfavorably for HUBOs by considering two test cases: a fully-connected four-local Hamiltonian, and the LABS problem. Next, we introduce in Section III an approximate quadratization procedure to reduce the QAOA circuit depth without additional qubits. Motivated by the idea of approximating the circuit Ansatz associated with a given problem, we benchmark how simplifications in the QAOA Ansatz improve measured approximation ratios when executing Max-Cut on hardware in Section IV. Finally, in Section V, we argue for the need of a more mathematically motivated quadratization and suggest some ideas in this direction.
@@ II Executing QAOA on digital quantum computers
We consider combinatorial optimization problems defined over $n$ binary variables $x\in\{0,1\}^{n}$ which minimize an objective function $f(x)$. Quantum computers can tackle such problems by mapping $f(x)$ to an Ising Hamiltonian $H_{C}$ whose ground state minimizes $f(x)$. This is done through the change of variables $x_{i}=\frac{1-z_{i}}{2}$ and promoting $z_{i}$ to Pauli $Z$ operators $Z_{i}$ [40]. Then, we use QAOA to find the ground state of $H_{C}$ by sampling from an Ansatz
|$\ket{\psi(\bm{\beta},\bm{\gamma})}=\prod_{q=1}^{p}e^{-i\beta_{q}H_{M}}e^{-i\gamma_{q}H_{C}}\ket{+}^{\otimes n}$ (1)
with optimized variational parameters $(\bm{\beta},\bm{\gamma})$. Here, we consider a problem-independent mixer operator $H_{M}=-\sum_{i=0}^{n-1}X_{i}$, although other mixing Hamiltonians have been developed to enforce constraints or leverage warm-starts [41, 42, 43, 28, 44, 45].
@@
scaling $n=16$ scaling $n=16$ $CZ$ gates $CZ$ gate depth # $CZ$ gates $CZ$ gate depth
$H_{2}^{\text{full}}$ $O(n^{2})$ 120 240 32 507 103
@@
QUBOs are unconstrained combinatorial optimization problems with a polynomial $f(x)$ of degree at most two. As such, their cost Hamiltonian is expressible as
$H_{\text{QUBO}}=\sum_{j<k}w_{jk}Z_{j}Z_{k}+\sum_{j}w_{j}Z_{j}.$ (2)
The main challenge associated with running QAOA on hardware is implementing the time evolution of the cost Hamiltonian. In particular, the more terms $H_{C}$ contains, the denser and deeper the resulting circuits become, which results in noisier samples. The maximum number of gates in the corresponding circuit of a fully-connected QUBO $H_{2}^{\text{full}}$ scales as $O(n^{2})$, with a $O(n)$ two-qubit gate depth [^1] We only consider two-qubit gates in gate counts as these are the main source of errors. For instance, the median error rates of the $CZ$ gates is $8.30\cdot 10^{-3}$, compared to $2.23\cdot 10^{-4}$ for single-qubit gates on ibm_fez. Similarly, when computing circuit depth, we refer only to the depth required to implement two-qubit gates., see the purple lines in Fig. 1. When transpiled to a line of qubits the complexity exponent stays the same, but the constant factor increases by 1.45 due to the additional SWAP gates required to reach all-to-all connectivity, see details in App. A and B.1.
Furthermore, the locality of the gates in $\exp(-i\gamma H_{C})$ significantly impacts the circuit depth. A gate $\exp(-i\gamma Z_{j}Z_{k})$ is implemented by a single $R_{ZZ}$ rotation if the qubits $j$ and $k$ are connected on the hardware. Otherwise, additional SWAP gates are needed to bring qubits $j$ and $k$ in contact. Minimizing the SWAP overhead is desirable since each SWAP gate is built from three noisy entangling gates. Implementing $\exp(-i\gamma H_{C})$ is even more challenging for HUBOs, whose terms require more stringent connectivity constraints, i.e. implementing $\exp(-i\gamma Z_{j}Z_{k}Z_{l}Z_{h})$ requires connecting qubits $(j,k,l,h)$. Therefore, HUBOs are more complicated to implement than QUBOs as they require implementing possibly many non-local gates. We now investigate the gate count and depth for the LABS problem and a fully-connected fourth-order Hamiltonian $H_{4}^{\text{full}}$.
The LABS problem has applications in signal-processing [46] and statistical physics [47]. It aims to minimize the sidelobe energy
@@
Expanding the sums in the expression of the sidelobe energy yields the Hamiltonian [37]
@@
Solving LABS with QAOA thus requires $O(n^{3})$ four-local terms and $O(n^{2})$ two-local terms. The dominant $O(n^{3})$ scaling is shown by the thicker blue lines in Fig. 1. For example, the largest LABS instance we solve has 16 decision variables, leading to 252 fourth-order terms and 56 second-order terms. This requires $252\times 6+56\times 2=1624$ two-qubits gates assuming an all-to-all connectivity, where the constant factors 6 and 2 are due to the two-qubit gate decomposition, see Figs. 2(a,b). By using the highest optimization level 3 when transpiling with Qiskit, this number reduces to 1218 through cancellation of neighboring two-qubit gates [48]. However, when compiled to linearly connected qubits, the two-qubit gate count increases to 4685 with a two-qubit gate depth of 2978, see the thin blue line in Fig. 1. Implementing such deep circuits on current noisy hardware is a formidable challenge. For instance, on ibm_fez, the median $T_{1}$ time of $146\leavevmode\nobreak\ {\rm\mu s}$ fits approximately only 1738 layers of two-qubit gates with a median gate time of $84\leavevmode\nobreak\ {\rm ns}$.
@@
The hardest fourth order problem to implement is a fully-connected four-local Hamiltonian $H_{4}^{\text{full}}$, which has a similar structure as the Sachdev-Kitaev-Ye model of Majorana fermions [49, 50, 51], but which is expressed in terms of Pauli operators instead of fermionic ones. Concretely, we define
$H_{4}^{\text{full}}=\sum_{j<k<h<l}^{{n\choose 4}}J_{jkhl}Z_{j}Z_{k}Z_{h}Z_{l}.$ (6)
@@
The weights $J_{jkhl}$ are sampled from the standard normal distribution. The resulting circuits have $O(n^{4})$ two-qubit gates, and $O(n^{3})$ two-qubit gate depth, see teal line in Fig. 1. Implementing QAOA for $H_{4}^{\text{full}}$ is thus more challenging than for LABS, and considerably more so than for a fully-connected QUBO. This is true not just in the asymptotic limit, but also for instances with as few as 16 variables which require 4732 $CZ$ gates and a depth of $4622$ assuming all-to-all connectivity, and more than three times as much for a linear qubit topology, see Tab. 1.
@@
Finally, these deep circuits result in low sampling rates on hardware. We estimate the rate at which we generate samples by the inverse of the duration of the quantum circuit. Furthermore, we lower bound the duration of a depth-one QAOA with the duration of running $e^{-i\gamma H_{C}}$, which is the product of the two-qubit gate duration $\bar{T}$ and the two-qubit gate depth. Extrapolating the data in Fig. 1 to a LABS system with 127 qubits, corresponding to one of the largest lines of superconducting qubits available [48], we conclude that we need a two-qubit gate depth of about $8.82\cdot 10^{6}$ to implement it. Taking $\bar{T}=84\leavevmode\nobreak\ {\rm ns}$, corresponding to the median $CZ$ gate duration of ibm_fez, this results in a runtime of $0.741\leavevmode\nobreak\ {\rm s}$ per sample. This means that we can generate samples at a rate of at most $1.35\leavevmode\nobreak\ {\rm s}^{-1}$, which is six orders of magnitude slower than the best-known classical methods for the same problem size [52].
@@ III Approximate quadratization
Compiling quartic HUBO problems to hardware-native instructions results in very deep circuits, while QUBOs are optimally compiled in linear depth with SWAP networks [41, 53]. This motivates us to design a QUBO based on the target HUBO to generate the Ansatz. For example, an exact quadratization that exactly conserves the full structure of the HUBO replaces each four-local term with one constant, four linear, and six quadratic terms at the expense of adding two ancilla qubits per four-local term [54]. This implies that implementing four-local Hamiltonians with $O(n^{3})$ terms, such as LABS, or $O(n^{4})$ terms, such as $H_{4}^{\text{full}}$, requires circuits with a qubit overhead scaling as $O(n^{3})$ or $O(n^{4})$, respectively. This is impractical due to the limited qubit count of the current hardware. We therefore introduce an approximate quadratization to project a generic HUBO onto a QUBO without adding decision variables.
III.1 Method

### v2

@@ Abstract
Combinatorial optimization problems have wide-ranging applications in industry and academia. Quantum computers may help solve them by sampling from appropriately prepared Ansatz quantum circuits. However, current quantum computers are limited by their qubit count, connectivity, and noise. This is particularly restrictive when considering optimization problems beyond the quadratic order. Here, we introduce Ansatze based on an approximate quadratization of high-order Hamiltonians which do not incur a qubit overhead. The price paid is a loss in the quality of the noiseless solution. Crucially, these approximations yield shallower Ansatze which are more robust to noise than the standard QAOA one. We show this through simulations with variable noise strengths. Furthermore, we also propose a noise-aware Ansatz design method for quadratic optimization problems. This method implements only a portion of the target Hamiltonian by limiting the number of layers of SWAP gates in the Ansatz. We find that for both problem types, under noise, our approximate implementation of the full problem structure can significantly enhance the solution quality. Our work opens a path to enhance the solution quality that approximate quantum optimization achieves on noisy hardware.
@@ I Introduction
Quantum computing offers a new information-processing framework based on the laws of quantum physics. For certain tasks, quantum computing achieves better performance than the best known classical methods. Famously, Shor’s algorithm delivers an exponential speed-up in factoring over state-of-the-art classical algorithms [shor1994algorithms]. However, this remains a theoretical feat since Shor’s algorithm requires deep circuits which cannot be faithfully implemented at practically relevant scales on current hardware due to noise. This has motivated the development of algorithms requiring shallower circuits such as Variational Quantum Algorithms (VQAs) [peruzzo2014variational]. VQAs are typically heuristic in nature and are often designed to find the ground state of a Hamiltonian [tilly2022variational, cao2018potential, blunt2022perspective]. VQAs optimize a parametrized quantum circuit, called the Ansatz, using a classical optimization procedure to update the parameters.
@@
The Ansatz in a VQA can be either problem-agnostic [kandala2017hardware] or retain the structure of the target problem [barkoutsos2018quantum]. Problem-agnostic Ansatze are typically designed to minimize the effect of noise by accounting for hardware constraints. However, they come at the cost of limited expressivity compared to Ansatze that fully encode the problem structure. Designing expressive and shallow Ansatze remains an open question [holmes2022connecting].
@@
Here, we focus on VQAs to solve combinatorial optimization problems, which are prevalent across industry applications such as communication [peterson2007computer], finance [markowitz2008portfolio, egger2020quantum], and vehicle routing [domino2022quadratic, pascariu2024formulation, sbihi2010combinatorial]. Many relevant combinatorial optimization problems are hard to solve exactly [karp1975computational]. In addition, for some of them, polynomial-time approximate algorithms that reach an arbitrarily accurate approximation of the optimal solution are not available [zuckerman1996unapproximable]. Therefore, many classical algorithms designed to solve such problems are heuristics. Crucially, they work well in practice, thus motivating a similar heuristic approach for quantum computing [abbas2024challenges].
@@
A foundational VQA for combinatorial optimization problems is the Quantum Approximate Optimization Algorithm (QAOA) [farhi2014quantum, blekos2024review]. Its trotterized Ansatz structure alternates $p$ time evolutions of a Hamiltonian representing the target problem and of a problem-independent Hamiltonian. The structure is inspired from the adiabatic theorem in physics, and thus inherits its convergence guarantee in the limit $p\to\infty$ [finnila1994quantum, farhi2014quantum, rajak2023quantum].
@@
Finite-depth QAOA is a heuristic algorithm, which implies that its performance must be thoroughly investigated and benchmarked on a case-by-case basis [abbas2024challenges, koch2025decathlon]. In particular, research often investigates QAOA for quadratic unconstrained binary optimization problems (QUBOs), such as Max-Cut, to benchmark hardware performance [zhou2020quantum, brandhofer2022benchmarking, willsch2020benchmarking, santra2024squeezing] and validate new algorithms [tate2023warm, egger2021warm]. However, Max-Cut is easy to solve classically at low graph densities and up to 10 000 nodes [rehfeldt2023faster]. Solving it using QAOA would only yield approximate solutions even in an ideal case [gamarnik2021overlap, lykov2023fast, lykov2023sampling, akshay2020reachability], thus making it a poor candidate to deliver a quantum advantage. Hence, recent studies investigated high-order unconstrained binary optimization problems (HUBOs) with cubic and quartic terms. Certain HUBOs can be very hard to solve, even at a low number of decision variables [romero2024bias]. For instance, the largest exactly solved instance of the Low Autocorrelation Binary Sequence (LABS) problem has only $66$ decision variables [brest2021low, packebusch2016low]. Importantly, Ref. [shaydulin2024evidence] suggests a scaling advantage for QAOA over Gurobi [gurobi] or CPLEX [cplex] in solving LABS.
@@
In this work, we first show in Section II how the quantum resources requirements of QAOA scale unfavorably for HUBOs by considering two test cases: a fully-connected four-local Hamiltonian, and the LABS problem. Next, we introduce in Section III two approximate quadratization procedures to reduce the QAOA circuit depth without qubit overhead. Motivated by the idea of approximating the circuit Ansatz associated with a given problem, we benchmark how simplifications in the QAOA Ansatz improve measured approximation ratios when executing Max-Cut on hardware in Section IV. Finally, in Section V, we argue for the need of a more mathematically motivated quadratization and suggest some ideas in this direction.
@@ II Executing QAOA on digital quantum computers
We consider combinatorial optimization problems defined over $n$ binary variables $x\in\{0,1\}^{n}$ which minimize an objective function $f(x)$. Quantum computers can tackle such problems by mapping $f(x)$ to an Ising Hamiltonian $H_{C}$ whose ground state minimizes $f(x)$. This is done through the change of variables $x_{i}=\frac{1-z_{i}}{2}$ and promoting $z_{i}$ to Pauli $Z$ operators $Z_{i}$ [lucas2014ising]. Then, we use QAOA to find the ground state of $H_{C}$ by sampling from an Ansatz
[…]
with optimized variational parameters $(\bm{\beta},\bm{\gamma})$. Here, we consider a problem-independent mixer operator $H_{M}=-\sum_{i=0}^{n-1}X_{i}$, although other mixing Hamiltonians have been developed to enforce constraints or leverage warm-starts [maciejewski2024design, zhu2022adaptive, fuchs2022constraint, egger2021warm, he2023alignment, niu2019optimizing].
@@
scaling $n=16$ scaling $n=16$ # $CZ$ gates CZCZ gate depth # $CZ$ gates CZCZ gate depth
$H_{2}^{\text{full}}$ - - $O(n^{2})$ 120 240 32 507 103
@@
QUBOs are unconstrained combinatorial optimization problems with a polynomial $f(x)$ of degree at most two. As such, their cost Hamiltonian is reads as
$H_{\text{QUBO}}=\sum_{i<j}w_{ij}Z_{i}Z_{j}+\sum_{j}w_{j}Z_{j}.$ (2)
The main challenge associated with executing QAOA on quantum processors is implementing the time evolution of the cost Hamiltonian. In particular, the more terms $H_{C}$ contains, the denser and deeper the resulting circuits become, which results in noisier samples. The maximum number of gates in the time-evolution circuit of a fully-connected QUBO $H_{2}^{\text{full}}$ scales as $O(n^{2})$, with a $O(n)$ two-qubit gate depth [^1] We only consider two-qubit gates in gate counts as these are the main source of errors. For instance, the median error rates of the $CZ$ gates is $8.30\cdot 10^{-3}$, compared to $2.23\cdot 10^{-4}$ for single-qubit gates on ibm_fez. Similarly, when computing circuit depth, we refer only to the depth required to implement two-qubit gates., see the purple lines in Fig. 1. When transpiled to a line of qubits the complexity exponent remains unchanged, but the scaling prefactor increases by 1.5 due to the additional SWAP gates required to reach all-to-all connectivity, see details in App. A.
Furthermore, the locality of the gates in $\exp(-i\gamma H_{C})$ significantly impacts the circuit depth. A gate $\exp(-i\gamma Z_{i}Z_{j})$ is implemented by a single $R_{ZZ}$ rotation if the qubits $i$ and $j$ are connected on the hardware. Otherwise, additional SWAP gates are needed to bring qubits $i$ and $j$ in contact. Minimizing the SWAP overhead is desirable since each SWAP gate is built from three noisy entangling gates. Implementing $\exp(-i\gamma H_{C})$ is even more challenging for HUBOs, whose terms require more stringent connectivity constraints, i.e. implementing $\exp(-i\gamma Z_{i}Z_{j}Z_{k}Z_{l})$ requires bringing close together qubits $(i,j,k,l)$. Therefore, HUBOs are more complicated to implement than QUBOs as they require implementing possibly many non-local gates. We now investigate the gate count and depth for two examples of HUBOs, i.e. the LABS problem and a fully-connected fourth-order Hamiltonian $H_{4}^{\text{full}}$.
The LABS problem has applications in signal-processing [golay977LABS] and statistical physics [bernasconi1987title]. It aims to minimize the sidelobe energy
@@
Expanding the sums in the expression of the sidelobe energy yields the Hamiltonian [shaydulin2024evidence]
@@
Solving LABS with QAOA requires implementing the time-evolution operator associated with $O(n^{3})$ four-local terms and $O(n^{2})$ two-local terms. The dominant $O(n^{3})$ scaling is shown by the thicker blue lines in Fig. 1. For example, a LABS instance with 16 decision variables has 252 four-local terms and 56 two-local terms. This requires $252\times 6+56\times 2=1624$ two-qubits gates assuming an all-to-all connectivity, where the constant factors 6 and 2 are due to the two-qubit gate decomposition, see Figs. 2(a,b). After transpilation using Qiskit with the highest optimization level 3, this number reduces to 1218 through cancellation of neighboring two-qubit gates [javadi2024quantum]. However, when compiled to linearly connected qubits, the two-qubit gate count increases to 4685 with a two-qubit gate depth of 2978, see the thin blue line in Fig. 1. Implementing such deep circuits on current noisy hardware is a formidable challenge. For instance, on ibm_fez, the median $T_{1}$ time of $146\penalty 10000\ {\rm\mu s}$ fits approximately only 1738 layers of two-qubit gates with a median gate time of $84\penalty 10000\ {\rm ns}$.
@@
The hardest fourth order problem to implement is a fully-connected four-local Hamiltonian $H_{4}^{\text{full}}$, which has a similar structure as the Sachdev-Kitaev-Ye model of Majorana fermions [rosenhaus2019introduction, luo2019quantum, kobrin2021many], but which is expressed in terms of Pauli operators instead of fermionic ones. Concretely, we define
$H_{4}^{\text{full}}=\sum_{i<j<k<l}^{{n\choose 4}}J_{ijkl}Z_{i}Z_{j}Z_{k}Z_{l},$ (6)
@@
$J_{ijkl}$ being real coefficients. The resulting circuits have $O(n^{4})$ two-qubit gates, and $O(n^{3})$ two-qubit gate depth, see teal line in Fig. 1. Implementing QAOA for $H_{4}^{\text{full}}$ is thus more challenging than for LABS, and considerably more so than for a fully-connected QUBO. This is true not only in the asymptotic limit, but also for instances with as few as 16 variables which require 4732 $CZ$ gates and a depth of $4622$ assuming all-to-all connectivity, and more than three times as much for a linear qubit topology, see Tab. 1.
@@
Finally, these deep circuits result in low sampling rates on hardware. We estimate the rate at which we generate samples by the inverse of the duration of the quantum circuit. Furthermore, we lower bound the duration of a depth-one QAOA with the duration of running $e^{-i\gamma H_{C}}$, which is the product of the two-qubit gate duration $\bar{T}$ and the two-qubit gate depth. Extrapolating the data in Fig. 1 to a LABS system with 127 qubits, corresponding to one of the largest lines of superconducting qubits available [javadi2024quantum], we conclude that it can be implemented with a two-qubit gate depth of about $8.82\cdot 10^{6}$. Taking $\bar{T}=84\penalty 10000\ {\rm ns}$, corresponding to the median $CZ$ gate duration of ibm_fez, this results in a runtime of $0.741\penalty 10000\ {\rm s}$ per sample. Therefore, we can generate samples at a rate of at most $1.35\penalty 10000\ {\rm s}^{-1}$, which is six orders of magnitude slower than the best-known classical methods for the same problem size [brest2018heuristic].
@@ III Approximate quadratization
Compiling quartic HUBO problems to hardware-native instructions results in very deep circuits, while QUBOs are optimally compiled in linear depth with SWAP networks [maciejewski2024design, matsuo2023sat]. This motivates us to generate the ansatz based on QUBOs defined in terms of the target HUBO. For example, an exact quadratization that exactly conserves the full structure of the HUBO replaces each four-local term with one constant, four linear, and six quadratic terms at the expense of adding two ancilla qubits per four-local term [mandal2020compressed]. This implies that implementing four-local Hamiltonians with $O(n^{3})$ terms, such as LABS, or $O(n^{4})$ terms, such as $H_{4}^{\text{full}}$, requires circuits with a qubit overhead scaling as $O(n^{3})$ or $O(n^{4})$, respectively. This is impractical due to the limited qubit count of the current hardware. We therefore explore approximate quadratization methods to project a generic HUBO onto a QUBO without adding decision variables.
III.1 Methods
We study two methods to quadratize HUBOs. The first is based on a hypergraph clique expansion, and the second is based on a fully-connected quadratic Hamiltonian whose gate weights are treated as additional variational parameters. The price paid to keep the qubit count constant is an approximation of the original problem.